# {tocentry} Vibrational Energy Dissipation in Non-Contact Single-Molecule Junctions Governed by Local Geometry and Electronic Structure

### v1

@@ 2.4 Electron-phonon coupling
Electron-phonon coupling is a significant factor in vibrational energy dissipation on metal surfaces.34, 2 Due to the absence of a band gap, the Born-Oppenheimer approximation does not strictly hold, and the motion of adsorbate atoms can induce EHP excitations that lead to nonadiabatic effects, sometimes also called electronic friction effects. In the weak perturbation regime, e.g. where adsorbate motion is triggered mechanochemically or through vibrational excitation, this effect leads to nonadiabatic energy dissipation. To describe the impact of these excitations on dynamics, different simulation methods have been developed. Time-dependent perturbation theory has emerged as a feasible approach for extended systems. This method enables the calculation of the electron-phonon-induced vibrational lifetimes (see Methods),35 which we discuss in the following.
@@
We evaluate electron-phonon coupling at a temperature of , typical for non-contact single-molecule junction experiments. At such low temperatures, electron-phonon coupling dominates over phonon-phonon coupling in metals.36, 37 At this temperature, first-order electron-phonon coupling is a dominant contribution to vibrational lifetimes and higher-order electron-phonon coupling effects can be neglected . 38 Its strength is sensitive to the electronic states that interact with nuclear motion, which we characterize by the electronic density of states (DOS) projected onto the copper atom bonded to CO (PDOS). The PDOS is highly sensitive to the geometry of the system. Since the energy to excite electrons is supplied by vibrational states with frequency , only a small portion of the DOS near the Fermi level ()—relative to the energy scale of electronic states—is accessible. Therefore, analysing the DOS at the Fermi level is instructive for understanding the strength of electron-phonon coupling. As shown in Figure 5, the PDOS of the sharp and blunt tips exhibits states localised in energy. The magnitude of the PDOS close to the Fermi level is small for the sharp tip, while the blunt tip exhibits energetically localised spikes. In contrast, the slab and slab-adatom systems have a PDOS that is delocalized in energy. The slab PDOS exhibits a larger magnitude compared to that of the adatom on the slab. Since a larger magnitude of the PDOS at and close to the Fermi level allows for stronger electron-phonon coupling, we expect that electron-phonon coupling will be stronger at flatter surfaces.
@@
The strength of electron-phonon coupling is reflected in the vibrational lifetimes, where shorter lifetimes indicate stronger coupling (see Table 4). The FT and SA modes experience relatively low electron-phonon coupling (resulting in longer lifetimes), whereas FR and IS modes exhibit electron-phonon coupling approximately an order of magnitude higher (with lifetimes an order of magnitude shorter) compared to FT and SA. Previous studies of CO on different Cu facets23, 22, 21 have reported similar trends of mode dependence, as we show in Table 3. Our results for the slab are in good agreement with literature values. In the case of the slab, there is an ongoing debate about whether the FR or the IS decays faster. Indeed, the lifetimes reported here and in other literature are very similar for these modes. However, as the tip becomes more corrugated, the IS mode clearly becomes the fastest dissipation mode, and the lifetime of the FR mode increases. As shown in Table 3, vibrational lifetimes are significantly longer on sharper tips, with CO on a flat surface exhibiting the shortest lifetimes. This trend aligns with the reduced magnitude of the PDOS close to the Fermi level on more corrugated surfaces.
@@ 2.5 Electron-phonon and phonon-phonon coupling
The experimentally measured lifetime is governed by both electron-phonon and phonon-phonon coupling. To capture this combined effect, we employ Langevin dynamics, specifically molecular dynamics with electronic friction (MDEF).29 We use a classical approximation, neglecting the zero-point energy of the phonons. This approach has been successfully applied to the dynamics of atoms39, 40 and molecules36, 37 at metal surfaces. In MDEF, electron-phonon coupling is treated as a friction force represented by the electronic friction tensor (see Equation 9). The electronic friction tensor is calculated in the Markovian quasi-static limit. Here, the frictional force experienced by the nuclei depends only on their instantaneous velocities, without memory (non-Markovian) effects from past states. This assumes that the electronic structure adjusts instantaneously to the nuclear configuration. In the quasi-static limit, the electronic friction tensor is computed for a given nuclear geometry without explicitly propagating electronic dynamics. This approach is common for molecular dynamics on metals.35, 41 We extract the total vibrational lifetime from phonon-phonon and electron-phonon contributions from MDEF simulations, by utilising two complementary methods. Equilibrium correlation analysis of the velocity cross-correlation function of MDEF trajectories in thermodynamic equilibrium allows determining lifetimes, as we have done for the phonon-phonon coupling. Kinetic energy decay predicts lifetimes by evaluating the explicit non-equilibrium decay of the kinetic energy of vibrational modes after selectively energizing a given mode. For both approaches, we employ normal mode decomposition to project the velocity time series onto the basis defined by the vibrational eigenvectors. Additionally, lifetime can be estimated from the independent phonon-phonon and electron-phonon lifetimes presented in Tables 3 and 2. Additive relaxation rates estimate the total lifetime by summing the phonon-phonon and electron-phonon relaxation rates predicted independently in previous sections.
@@
Equilibrium correlation analysis. Table 4 shows the vibrational lifetimes from equilibrium correlation analysis of the MDEF simulation results. While the IS and FR lifetimes are dominated by the electron-phonon contribution, the FT mode is dominated by the phonon-phonon contribution. The two effects are somewhat equal in contribution to the SA mode lifetime. Lifetimes for the FR and IS modes are in good agreement with infrared spectroscopy experiment. However, the experimental lifetime of the SA mode is , while the calculated lifetime is for the Cu(111) surface. A similar discrepancy has been reported in literature and was attributed to inhomogeneous broadening:6 Experiments conducted on rougher surfaces show a greater sensitivity of the mode amplitude and linewidth of the SA mode (and the amplitude of the FR mode) than for the IS mode. We note in passing that the literature lifetimes from infrared spectroscopy experiments should be considered as lower bounds (marked by asterisks (*) in Table 4). This is because the observed linewidths are artificially broadened due to limitations of the measurement setup. Since lifetime is inversely related to linewidth, this broadening leads to a systematic underestimation of the lifetimes. Additionally, helium scattering experiments have measured the lifetime of the FT mode of CO on Cu(100) to ,42 which is in the same order of magnitude as the lifetime for the FT mode of CO on a Cu(111) slab reported here.
@@
Kinetic energy decay. We can compare lifetimes determined through equilibrium correlation analysis with those explicitly calculated from non-equilibrium simulations (see Method section). In these simulations, we excite vibrations with a specific amount of energy and monitor the decay of the kinetic energy (Figure 6). To determine the lifetimes, we assume that the vibrational modes behave like damped harmonic oscillators, where the envelope of the kinetic energy is assumed to exhibit an exponential decay. Overall, we observe that the kinetic energies of the CO vibrational modes indeed show an exponential decay. A comparison between Table 4 and Figure 6 demonstrates that the non-equilibrium lifetimes agree well with those from equilibrium correlation analysis.
@@
Additive relaxation rates. The total relaxation rate , or the inverse of the lifetime , can be approximated by summing the contributions from electron-phonon (Table 3) and phonon-phonon coupling (Table 2).
@@
Figure 7a compares the three methods of determining the total lifetime (or relaxation rate). Within the estimated uncertainty of our simulations, the results are in good agreement. All three methods yield lifetimes that exhibit similar trends across the different systems and modes. Flatter surface geometries lead to overall larger relaxation rates and smaller lifetimes. Relaxation rates exhibit a strong mode dependence, with the largest rate being an order of magnitude greater than the smallest. The numerical values in Figure 7a can be found in Section S3 of the Supporting Information.
@@
Contributions to the relaxation rates. Let us now disentangle the phonon-phonon and electron-phonon contributions to the relaxation rates in dynamic simulations. Figure 7b shows the two components that we decompose using the kinetic energy decay method. The phonon-phonon relaxation rates are determined using a similar procedure to that for the total relaxation rate, but exclusively with MD and without electronic friction. The phonon-phonon relaxation rates are obtained by monitoring the decay in kinetic energy during MD simulations (similar to total rates but without electronic friction). The electron-phonon relaxation rate is extracted from the MDEF trajectories by integrating the electronic friction and random force components (see equation 9). We note that the random force component integrates to zero in equilibrium.
@@
Figure 7b shows that overall, the relaxation rate (and consequently the lifetimes) of the FT and SA modes are primarily dominated by phonon-phonon coupling, whereas the FR and IS modes are predominantly influenced by electron-phonon coupling. Interestingly, the sum of phonon-phonon and electron-phonon contributions does not exactly match the total relaxation rate. Overall, we observe a trend to underestimate the total rate if individual components are summed. While this trend to underestimate the total relaxation rate is within the uncertainty of our simulation, it does hint towards some level of non-additivity and a deviation from equation 5 in the MDEF dynamics. Moreover, we have assumed a constant electronic friction tensor. In reality, the friction tensor is configuration dependent and changes as the geometry is displaced along different vibration modes. This effect will likely further add to non-additivity. Electron-phonon and phonon-phonon relaxation rates may not be fully independent, as shown in Figures S18, S19, S20, and S21 in the Supporting Information. There, we find that the phonon-phonon coupling between the CO and the substrate vibrational modes increases if electron-phonon coupling is included in the equilibrium dynamics simulations.
@@ Discussion
Vibrational modes and phonon-phonon coupling. Remarkably, all phononic lifetimes fall within the same order of magnitude. In contrast, electron-phonon lifetimes vary by two orders of magnitude (IS vs. FT), highlighting a much stronger mode dependence compared to phonon-phonon interactions. Additionally, despite the significantly higher vibrational energy of the IS mode, its lifetime remains comparable to other modes, suggesting a surprising degree of energy independence in phonon-phonon coupling. To understand this behaviour, we analyze the VDOS and coupling density. In the slab and slab-adatom systems, energetically more localized substrate vibrational states lead to reduced phonon-phonon coupling compared to the sharp and blunt tips. This is counteracted by a larger number of vibrational states per volume (approximately three times as many) in the slab and slab-adatom systems. The FT mode has an energy comparable to the lowest substrate vibrational modes, while the FR mode is close in energy to the largest substrate modes. Coupling of the molecular vibrations is most efficient to the substrate vibrations with similar energy, as shown in Figure 4 and Supporting Information Figures S18, S19, S20, and S21 (Section S5). We can understand this in the following way: Energy conservation mandates that any newly excited vibrational mode arises from the creation or annihilation of other modes, following . As the number of participating modes increases, the efficiency of the process decreases. Because the FT and FR modes overlap energetically with the substrate modes, they can access efficient processes. However, individual substrate vibrations are strongly dependent on the surface shape. As a result, the phononic lifetimes of the FT and FR modes depend strongly on the surface shape, as shown in Table 2. The SA and IS modes are energetically relatively decoupled from the substrate modes compared to FT and FR. Therefore, their coupling of the substrate modes involves high- processes. Supporting Information Figures S20 and S21 show that the SA and IS modes couple equally to all surface modes. Consequently, the influence of surface geometry on specific surface vibrations is effectively averaged out, as shown in Table 2.
@@
Electron-phonon coupling. We find small electron-phonon lifetimes for the FR and IS modes and large lifetimes for the FT and SA modes, in good agreement with experimental and theoretical literature.6, 21, 23, 22 For each tip geometry, the electron-phonon lifetimes follow the trend of the vibrational frequencies of the respective modes, except for the FR mode. We can understand this behavior if we assume that the rate of EHP excitation is independent of the energies of the initial and final states. This corresponds to the wide-band, constant-coupling approximation, where one assumes a constant electronic DOS. In such a case, relaxation rates scale quadratically with the vibration energy. A similar explanation has been proposed by Persson.23 As the vibrational energy increases, the available energy window for electronic transitions broadens, enabling more excitation pathways and thereby decreasing the lifetime. The vibrational energy of the IS mode is two orders of magnitude higher than that of the FT mode, consistent with the two-order-of-magnitude difference in their lifetimes. The FR mode deviates from this trend, with lifetimes comparable to the IS mode for the slab geometry. This can be explained by the large electron-phonon coupling reported by STM vibrational spectroscopy experiments for this mode.44 Therefore, the EHP rate cannot be assumed to be independent of the involved energy states in this case.
@@
Moving from the slab model to the sharp tip model, the lifetimes become larger. As Figure 5 shows, the magnitude of the electronic PDOS becomes smaller with more corrugated tips. This leaves fewer electronic states available for electronic excitations. Consequently, electron-phonon coupling becomes less efficient, and the lifetime increases. However, the IS mode deviates from this trend. The available vibrational energy of the IS mode is approximately , which is one order of magnitude larger than the energy of the SA and FR modes. The large vibrational energy of the IS mode increases the number of accessible electronic states for excitation. Even for the sharp tip, where the PDOS is highly localized, this energy range enables excitations between many more electronic states. Since the other modes carry significantly less energy, they are more strongly affected by the surface geometry: The lifetime of the FT mode increases by one order of magnitude between the slab and the sharp tip, making it two orders of magnitude longer than that of the IS mode.
@@
Combined energy dissipation mechanisms. We have determined total relaxation rates using three methods: additive relaxation rates, equilibrium correlation analysis, and kinetic energy decay. Despite their significant conceptual difference, all three methods predict lifetimes that exhibit similar trends across the different systems and modes (see Figure 7a). This demonstrates the robustness of our analysis.
@@
In Figure 7b, we have also observed that the sum of phonon-phonon and electron-phonon relaxation rates from dynamic simulations tends to be smaller than the MDEF relaxation rates. While this difference lies within the uncertainties and is therefore strictly not significant, it is in line with a comparison of the coupling densities with and without electron-phonon coupling: Coupling densities become larger if electron-phonon coupling is included in the simulations, indicating a weak non-additive effect where electron-phonon-mediated phonon-phonon coupling increases phononic relaxation. However, the approximations in our approach likely underestimate this non-additivity. We assume a constant electronic friction tensor, whereas in reality, it is configuration-dependent. As the geometry is displaced along different vibration modes, the strength of electron-phonon coupling changes, likely enhancing the degree of non-additivity.
@@
The observation from Figure 7b that the relaxation rate of the FR (and IS) mode is dominated by electron-phonon coupling while the SA mode is not, is in good agreement with experiment. Asymmetric Fano line shapes have been reported for the FR mode, while symmetric Lorentzian line shapes have been measured for the SA mode.6 Fano line shapes typically indicate that a discrete vibrational mode interacts with a continuum of electronic states, indicating significant electron-phonon coupling. Conversely, a Lorentzian line shape suggests that phonon-phonon coupling dominates the relaxation of the vibration mode.45
@@
However, it is important to note that the presence of electron-phonon coupling does not necessarily lead to a Fano line shape. According to Langreth,45 the line shape of a vibrational mode in the presence of EHP excitations is governed by the dynamic dipole , with the asymmetry depending on the imaginary part, :
@@
Here and are the ionic and electronic dipole, respectively. The electronic dipole consists of a real and an imaginary part. The imaginary part emerges due to the creation of EHPs as adsorbate electrons and holes tunnel into the substrate. Since this tunneling process is not instantaneous, the oscillations of the electronic dipole get out of phase with the driving electric field.
@@
The imaginary part of equation 6 becomes large—giving rise to an asymmetric line shape—if the imaginary electronic dipole is large compared to the ionic dipole . This occurs for modes with a dipole oriented parallel to the substrate—such as the FR mode—where screening by mirror charges significantly reduces the effective dipole strength. By contrast, modes with atomic motion perpendicular to the substrate, such as the SA and IS modes, exhibit a large ionic dipole. Subsequently, asymmetric line shapes would not be expected here, even if electron-phonon coupling is the dominant mechanism.

### v2

@@ 2.4 Electron-phonon coupling
EPC is a significant factor in vibrational energy dissipation on metal surfaces.37, 2 Due to the absence of a band gap, the Born-Oppenheimer approximation does not strictly hold, and the motion of adsorbate atoms can induce EHP excitations that lead to nonadiabatic effects, sometimes also called electronic friction effects. In the weak perturbation regime, e.g. where adsorbate motion is triggered mechanochemically or through vibrational excitation, this effect leads to nonadiabatic energy dissipation. To describe the impact of these excitations on dynamics, different simulation methods have been developed. Time-dependent perturbation theory has emerged as a feasible approach for extended systems. This method enables the calculation of the electron-phonon-induced vibrational lifetimes (see Description of electron-phonon coupling in the Methods),38 which we discuss in the following.
@@
We evaluate EPC at a temperature of , typical for non-contact single-molecule junction experiments. At such low temperatures, EPC dominates over PPC in metals 39, 40 and first-order EPC is the dominant contribution to vibrational lifetimes, allowing higher-order EPC effects to be neglected.41 Its strength is sensitive to the electronic states that interact with nuclear motion, which we characterize by the electronic density of states (DOS) projected onto the copper atom bonded to CO (PDOS), which is obtained from Density Functional Theory (DFT). The PDOS is highly sensitive to the geometry of the system. Since the energy to excite electrons is supplied by vibrational states with frequency , only a small portion of the DOS near the Fermi level ()—relative to the energy scale of electronic states—is accessible. Therefore, analysing the DOS at the Fermi level is instructive for understanding the strength of EPC. As shown in Figure 5, the PDOS of the sharp and blunt tips exhibits states localised in energy. The magnitude of the PDOS close to the Fermi level is small for the sharp tip, while the blunt tip exhibits energetically localised spikes. In contrast, the slab and slab-adatom systems have a PDOS that is delocalized in energy. The slab PDOS exhibits a larger magnitude compared to that of the adatom on the slab. Since a larger magnitude of the PDOS at and close to the Fermi level allows for stronger EPC, we expect that EPC will be stronger at flatter surfaces.
@@
The strength of EPC is reflected in the vibrational lifetimes, where shorter lifetimes indicate stronger coupling (see Table 3). The FT and SA modes experience relatively low EPC (resulting in longer lifetimes), whereas FR and IS modes exhibit EPC approximately an order of magnitude higher (with lifetimes an order of magnitude shorter) compared to FT and SA. Previous studies of CO on different Cu facets25, 24, 23 have reported similar trends of mode dependence, as we show in Table 3. Our results for the slab are in good agreement with literature values. In the case of the slab, there is an ongoing debate about whether the FR or the IS decays faster. Indeed, the lifetimes reported here and in other literature are very similar for these modes. However, as the tip becomes more corrugated, the IS mode clearly becomes the fastest dissipation mode, while the lifetime of the FR mode increases. As shown in Table 3, vibrational lifetimes are significantly longer on sharper tips, with CO on a flat surface exhibiting the shortest lifetimes. This trend aligns with the reduced magnitude of the PDOS close to the Fermi level on more corrugated surfaces.
@@ 2.5 Electron-phonon and phonon-phonon coupling
The experimentally measured lifetime is governed by both EPC and PPC. To capture this combined effect, we employ Langevin dynamics, specifically MDEF.31 We use a classical approximation, neglecting the zero-point energy of the phonons. This approach has been successfully applied to the dynamics of atoms42, 43 and molecules39, 40 at metal surfaces. In MDEF, EPC is treated as a friction force represented by the electronic friction tensor (see Equation 9). The electronic friction tensor is calculated in the Markovian quasi-static limit (see Electron-phonon coupling). Here, the frictional force experienced by the nuclei depends only on their instantaneous velocities, and memory effects are ignored. This assumes that the electronic structure adjusts instantaneously to the nuclear configuration. In the quasi-static limit, the electronic friction tensor is computed at zero frequency in the harmonic approximation for a given nuclear geometry without explicitly propagating electronic dynamics. This approach is commonly employed for MDEF on metals.38, 44 We determine vibrational relaxation rates and lifetimes using three different methods (see Figure 6a):
Equilibrium correlation analysis. Lifetimes are determined by analyzing the velocity cross-correlation function from an MD (without EPC) or MDEF (with EPC) trajectory in thermodynamic equilibrium. This method is introduced and applied in Phonon-phonon coupling to determine PPC relaxation rates and lifetimes from MD trajectories. When EPC is included via MDEF, the method yields total relaxation rates that capture both PPC and EPC.
@@
Table 4 shows the vibrational lifetimes from equilibrium correlation analysis of the MDEF simulation results. While the IS and FR lifetimes are dominated by EPC, the FT mode is dominated by PPC. The two effects are somewhat equal in contribution in the case of the SA mode lifetime. Lifetimes for the FR and IS modes are in good agreement with infrared spectroscopy experiment. However, the experimental lifetime of the SA mode is , while the calculated lifetime is for the Cu(111) surface. A similar discrepancy has been reported in literature and was attributed to inhomogeneous broadening:6 Experiments conducted on rougher surfaces show a greater sensitivity of the mode amplitude and linewidth of the SA mode (and the amplitude of the FR mode) than for the IS mode. The literature lifetimes from infrared spectroscopy experiments should be considered as lower bounds (marked by asterisks (*) in Table 4). This is because the observed linewidths are additionally broadened due to limitations of the measurement setup. Since lifetime is inversely related to linewidth, this broadening leads to a systematic underestimation of the lifetimes. Additionally, helium scattering experiments have measured the lifetime of the FT mode of CO on Cu(100) to ,45 which is in the same order of magnitude as the lifetime for the FT mode of CO on a Cu(111) slab reported here.
@@
Kinetic energy decay. We can compare lifetimes determined through equilibrium correlation analysis with those explicitly calculated from non-equilibrium kinetic energy decay simulations (see Method section). Hereby, lifetimes are predicted by monitoring how the kinetic energy of an initially excited vibration mode decreases over time during a non-equilibrium MDEF simulation. We excite vibrations with a specific amount of energy and monitor the decay of the kinetic energy (Figure S18 in the Supporting Information). To determine the lifetimes, we assume that the vibrational modes behave like damped harmonic oscillators, where the envelope of the kinetic energy is assumed to exhibit an exponential decay. By including EPC using MDEF, total relaxation rates can be obtained. Overall, we observe that the kinetic energies of the CO vibrational modes indeed show an exponential decay (see Figure S18 in the Supporting Information).
@@
Additive relaxation rates. We sum two independent contributions to the vibrational relaxation rate: the PPC relaxation rates obtained from MD equilibrium correlation analysis, and the EPC relaxation rates calculated in the quasi-static limit using first-order time-dependent perturbation theory. The EPC contribution calculated in this way is described in detail in Electron-phonon coupling. The total relaxation rate is the inverse of the lifetime and can be approximated by summing the contributions from PPC (Table 2) and EPC (Table 3):
@@
Figure 6b compares the three methods of determining total lifetimes (and relaxation rates). Additional technical details on the three methods are provided in Section S3 of the Supporting Information. Within the estimated uncertainty of our simulations, the results are in good agreement. All three methods yield lifetimes that exhibit similar trends across the different systems and modes. CO adsorption on a terrace leads to overall larger relaxation rates and smaller lifetimes compared to adsorption on adatoms or single atom tips. Relaxation rates exhibit a strong mode dependence, with the largest rate being an order of magnitude greater than the smallest. The numerical values in Figure 6b can be found in Section S3 of the Supporting Information.
@@
Contributions to the relaxation rates. We decompose the total vibrational relaxation rate into EPC and EPP contributions to identify the dominant dissipation channel for each vibrational mode in each adsorption environment. This is done using the kinetic energy decay method. The phonon-phonon relaxation rates are obtained by monitoring the decay in kinetic energy during MD simulations (which do not include EPC). The electron-phonon relaxation rate is extracted from the MDEF trajectories. This is accomplished by integrating the electronic friction and random force components in equation 9 along the trajectory.
Results for all adsorption environments are presented in Figure S22 in the Supporting Information and specifically for the slab adsorption environment in Figure 6c. The relative contributions of PPC and EPC to the total relaxation rate are relatively independent of the adsorption environment as shown in Figure S22 in the Supporting Information, but depend strongly on the vibrational mode. PPC primarily dominates the relaxation of FT and SA modes, whereas EPC predominantly influences the FR and IS modes.
@@
Interestingly, the contributions from EPC and PPC do not fully add up to the total relaxation rate (see non-additive component in Figure 6c) even when considering statistical uncertainties in the analysis. We observe a trend to underestimate the total rate, and the non-additivity is more pronounced in modes where EPC dominates (FR and IS modes). EPC and PPC may not be fully independent, as shown in Figures S24, S25, S26, and S27 in the Supporting Information. There, we find that PPC between the CO and the substrate vibrational modes increases if EPC is included in the equilibrium dynamics simulations. This result hints towards some level of non-additivity and a deviation from equation 5 when considering both effects simultaneously in the MDEF dynamics. Moreover, certain approximations taken in this work are expected to lead to more additivity. For example, we have assumed a configuration-independent electronic friction tensor. In reality, the friction tensor is configuration dependent and changes as the geometry is displaced along different vibration modes, likely contributing further to non-additivity.
@@ Discussion
Vibrational modes and phonon-phonon coupling. Remarkably, all phononic lifetimes fall within the same order of magnitude. In contrast, electron-phonon lifetimes vary by two orders of magnitude (IS vs. FT), highlighting a much stronger mode dependence compared to phonon-phonon interactions. Additionally, despite the significantly higher vibrational energy of the IS mode, its lifetime remains comparable to other modes, suggesting a surprising degree of energy independence in PPC. To understand this behaviour, we analyze the VDOS and coupling density. In the slab and slab-adatom systems, energetically more localized substrate vibrational states lead to reduced PPC compared to the sharp and blunt tips. This is counteracted by a larger number of vibrational states per volume (approximately three times as many) in the slab and slab-adatom systems. The FT mode has an energy comparable to the lowest-energy substrate vibrational modes, while the FR mode is close in energy to the highest-energy substrate modes. Coupling of the molecular vibrations is most efficient to the substrate vibrations with similar energy, as shown in Figure 4 and Supporting Information Figures S24, S25, S26, and S27 (Section S6.3). We can understand this in the following way: Energy conservation mandates that any newly excited vibrational mode arises from the creation or annihilation of other modes, following . As the number of participating modes increases, the efficiency of the process decreases. Because the FT and FR modes overlap energetically with the substrate modes, they can access efficient processes. However, individual substrate vibrations are strongly dependent on the surface shape. As a result, the phononic lifetimes of the FT and FR modes depend strongly on the surface shape, as shown in Table 2. The SA and IS modes are energetically relatively decoupled from the substrate modes compared to FT and FR. Therefore, their coupling of the substrate modes involves high- processes. Supporting Information Figures S26 and S27 show that the SA and IS modes couple equally to all surface modes. Consequently, the influence of surface geometry on specific surface vibrations is effectively averaged out, as shown in Table 2.
@@
Electron-phonon coupling. We find small electron-phonon lifetimes for the FR and IS modes and large lifetimes for the FT and SA modes, in good agreement with experimental and theoretical literature.6, 23, 25, 24 For each adsorption environment, the electron-phonon lifetimes follow the trend of the vibrational frequencies of the respective modes, except for the FR mode. We can understand this behavior if we assume that the rate of EHP excitation is independent of the energies of the initial and final states. This corresponds to the wide-band, constant-coupling approximation, where one assumes a constant electronic DOS. In such a case, relaxation rates scale quadratically with the vibration energy. A similar explanation has been proposed by Persson.25 As the vibrational energy increases, the available energy window for electronic transitions broadens, enabling more excitation pathways and thereby decreasing the lifetime. The vibrational energy of the IS mode is two orders of magnitude higher than that of the FT mode, consistent with the two-order-of-magnitude difference in their lifetimes. The FR mode deviates from this trend, with lifetimes comparable to the IS mode for the slab geometry. This can be explained by the large EPC reported by STM vibrational spectroscopy experiments for this mode.47 Therefore, the EHP rate cannot be assumed to be independent of the involved energy states in this case.
@@
Moving from the slab model to the sharp tip model, the lifetimes become larger. As Figure 5 shows, the magnitude of the electronic PDOS becomes smaller with more corrugated adsorption environments. This leaves fewer electronic states available for electronic excitations. Consequently, EPC becomes less efficient, and the lifetime increases. However, the IS mode deviates from this trend. The available vibrational energy of the IS mode is approximately , which is one order of magnitude larger than the energy of the SA and FR modes. The large vibrational energy of the IS mode increases the number of accessible electronic states for excitation. Even for the sharp tip, where the PDOS is highly localized, this energy range enables excitations between many more electronic states. Since the other modes carry significantly less energy, they are more strongly affected by the adsorption environment: The lifetime of the FT mode increases by one order of magnitude between the slab and the sharp tip, making it two orders of magnitude longer than that of the IS mode.
@@
Combined energy dissipation mechanisms. We have determined total relaxation rates using three methods: equilibrium correlation analysis, kinetic energy decay, and additive relaxation rates. Despite their significant conceptual difference, all three methods predict lifetimes that exhibit similar trends across the different systems and modes (see Figure 6b). This demonstrates the robustness of our analysis.
@@
In Figure 6c, the sum of phonon-phonon relaxation rates from MD and electron-phonon rates from MDEF simulations tends to be smaller than the total relaxation rates obtained directly from MDEF. For most modes, this difference lies within the uncertainties and is therefore strictly not statistically significant; it is in line with a comparison of the coupling densities with and without EPC: Coupling densities become larger if EPC is included in the simulations, indicating a weak non-additive effect where electron-phonon-mediated PPC increases phononic relaxation. However, the approximations in our approach likely underestimate this non-additivity. We assume a constant electronic friction tensor, whereas in reality, it is configuration-dependent. As the geometry is displaced along different vibration modes, the strength of EPC changes, likely enhancing the degree of non-additivity.
@@
The observation from Figure 6c that the relaxation rate of the FR (and IS) mode is dominated by EPC while the SA mode is not, is in good agreement with experiment. Asymmetric Fano line shapes have been reported for the FR mode, while symmetric Lorentzian line shapes have been measured for the SA mode.6 Fano line shapes typically indicate that a discrete vibrational mode interacts with a continuum of electronic states, indicating significant EPC. Conversely, a Lorentzian line shape suggests that PPC dominates the relaxation of the vibration mode.48
@@
However, it is important to note that the presence of EPC does not necessarily lead to a Fano line shape. Following Langreth,48 the line shape of a vibrational mode in the presence of EHP excitations is governed by the dynamic dipole , with the asymmetry depending on the imaginary part of the equation:
@@
Here and are the ionic and electronic dipole, respectively. The electronic dipole consists of a real and an imaginary part. The imaginary part, , emerges due to the creation of EHPs as adsorbate electrons and holes tunnel into the substrate. Since this tunneling process is not instantaneous, the oscillations of the electronic dipole get out of phase with the driving electric field.
@@
The imaginary part of equation 6 becomes large—giving rise to an asymmetric line shape—if the imaginary electronic dipole is large compared to the ionic dipole . This occurs for modes with a dipole oriented parallel to the substrate—such as the FR mode—where screening by mirror charges significantly reduces the effective dipole strength. By contrast, modes with atomic motion perpendicular to the substrate, such as the SA and IS modes, exhibit a large ionic dipole. Subsequently, asymmetric line shapes would not be expected here, even if EPC is the dominant mechanism.
Our findings also inform the expected dynamic properties of CO on other coinage metals, as discussed in Section S4 of the Supporting Information. The greater mass mismatch between CO and Ag or Au, compared to Cu, reduces PPC. Moreover, the phonon density of states shifts to lower energies from Cu to Ag to Au (Figure S23 in the Supporting Information), leading to greater decoupling of the IS mode on Ag and Au than on Cu. Given the chemical similarity of the coinage metals, all of which are predominantly s-band conductors, this trend is expected to persist for different adsorption environments. Assuming comparable adsorption conditions, CO exhibits similar vibrational lifetimes on Cu, Ag, and Au.49, 50 However, due to weaker interactions with Ag and Au, EPC is significantly reduced.49 For adatoms or tip-like adsorption environments, trends observed on Cu likely extend to Ag and Au, although stronger interactions at low-coordination sites may facilitate shorter lifetimes relative to flat surfaces.
Perspective and potential experimental validation. Finally, vibrational lifetimes calculated with equilibrium correlation analysis and kinetic energy decay are in good agreement, which demonstrates that both are suitable and equivalent approaches. Equilibrium correlation analysis can yield important information, such as coupling densities and phononic spectral functions. Our method for disentangling electron-phonon and phonon-phonon contributions to the vibrational dynamics of adsorbed molecules enables a better understanding of electronic and vibrational mass effects, as well as effects due to the local bonding environment. To connect theory with experiment, it is essential to identify techniques capable of validating mode-specific dissipation pathways. Most existing experiments for systems such as CO on Cu were conducted decades ago and are limited to flat surfaces.6, 7 Data is lacking on low-energy modes, such as the FT. High resolution IR (infrared) and Raman spectroscopy allow for directly probing the dominant dissipation mechanisms via vibrational lineshapes: for vibrational modes with a small ionic dipole (such as those parallel to the surface where the dipole is screened), dominant EPC leads to characteristic Fano profiles (see Equation 6), while PPC results in Lorentzian lineshapes.48
Tip-enhanced Raman spectroscopy (TERS) offers a powerful way to probe vibrational dynamics at the single-molecule level with atomic resolution, particularly when operated in the quantum tunneling regime of plasmonic enhancement.51 By selectively tuning the excitation laser to match vibronic transitions, it becomes possible to achieve highly mode-specific resonance enhancement.52 In STM-TERS, the application of a tunable bias voltage provides additional control: it can alter the tunneling gap and modulate the strength of EPC, which can be tracked via bias-dependent frequency shifts and intensity changes.53 This could help to disentangle phonon-phonon and electron-phonon relaxation pathways and their contributions at specific adsorption sites.
To account for the adsorption site dependence (CO on different adsorption environments), experiments with spatial resolution are needed. STM-IETS (inelastic electron tunneling spectroscopy) can resolve low energy modes, is sensitive to IR and Raman inactive modes, and enables atomic spatial resolution.3 Crucially, it is highly sensitive to EPC: Electrons lose energy by exciting vibrations during tunneling, and the resulting peak intensity scales with the electron–phonon coupling strength.54, 55 Meanwhile, the peaks are broadened by both EPC and PPC. By combining IETS with a different method of vibrational excitation, such as IR, which excites modes via changes in the dipole moment, it becomes possible to disentangle the dissipation mechanisms. For instance, a mode showing a large IR peak but a small IETS signal would indicate that PPC dominates, while a strong IETS response would indicate that EPC is significant.